OpenStudy (anonymous):

The number of milligrams of a drug that remains in a patient's system after t hours is given by the function A(t) = Iert. Juan was given 500 milligrams of medicine which leaves his bloodstream at a rate of 25%. How much of the medicine remains in his system after 6 hours?

OpenStudy (anonymous):

A ( 6) = 500 * e^ ( -.25 * 6) Remain amount: 500 = A (6 ) = ...?

OpenStudy (anonymous):

@kerber123 Can you do it?

OpenStudy (anonymous):

A=111.57 ?

OpenStudy (anonymous):

@Chlorophyll thank you!

OpenStudy (anonymous):

Do you have the answer? ( I don't have the calculator with me )

OpenStudy (anonymous):

Find the remain!